OpenStudy (anonymous):

how do I issolate 'y' in this equasion. . . . xy-2y^2=-4

jimthompson5910 (jim_thompson5910):

\[\large xy-2y^2=-4\] \[\large xy-2y^2+4 = 0\] \[\large -2y^2+xy+4 = 0\] The last equation is in the form ay^2 + by + c = 0 where a = -2 b = x c = 4 From here, you plug all this into the quadratic formula given below, and simplify, to solve for y. \[\large y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] I'll let you take over from here.

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

this is what i got \[y=\frac{ x \pm \sqrt{x^2+32} }{ -4 }\] what do i do about the plus or minus?

jimthompson5910 (jim_thompson5910):

Here's how I would do it \[\large y=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\large y=\frac{-x \pm \sqrt{(-x)^2-4(-2)(4)}}{2(-2)}\] \[\large y=\frac{-x \pm \sqrt{x^2+32}}{-4}\] \[\large y=\frac{-x + \sqrt{x^2+32}}{-4} \ \text{or} \ \frac{-x - \sqrt{x^2+32}}{-4}\] \[\large y=\frac{x - \sqrt{x^2+32}}{4} \ \text{or} \ \frac{x + \sqrt{x^2+32}}{4}\] \[\large y=\frac{x + \sqrt{x^2+32}}{4} \ \text{or} \ \frac{x - \sqrt{x^2+32}}{4}\] So you just needed to break up the +- into two different equations

OpenStudy (anonymous):

ok. that makes sense

jimthompson5910 (jim_thompson5910):

ok great, glad it does